OpenStudy (anonymous):

Which of the following is the solution set of 18 - 3n + 2 = n + 20 - 4n?

OpenStudy (solomonzelman):

MEDAL FOR THE ANSWER?

OpenStudy (anonymous):

Yeahhhh (:

OpenStudy (solomonzelman):

The equation is 18 - 3n + 2 = n + 20 - 4n -3n + 4n - n = 20 -18 - 2 0 = 0 for n = 0 18 - 3*0 + 2 = 0 + 20 - 4*0 18 + 2 = 20 20 = 20 SO for n = 1 18 - 3*1 + 2 = 1 + 20 - 4*1 18 - 3 + 2 = 1 + 20 - 4 17 = 17 for n = -1 18 - 3*(-1) + 2 = -1 + 20 - 4*(-1) 18 + 3 + 2 = -1 + 20 + 4 23 = 23 for n = -2 18 -3*(-2) + 2 = -2 + 20 -4*(-2) 18 + 6 + 2 = -2 + 20 + 8 26 = 26 so n = [-infinity , ..., 0 , ... , infinity]

OpenStudy (solomonzelman):

GOOD?

OpenStudy (anonymous):

Yes .. Thankyouuu (: ....

OpenStudy (solomonzelman):

Bye!

OpenStudy (anonymous):

Byeee (: